# Stable High-Order Interpolation on the Grassmann Manifold by Maximum-Volume Coordinates and Arnoldi Orthogonalization

### 25 Mar 2026 v1

@@ 5.2 Algebraic Stability of CV+A
Let $p(t)=\sum_{j=0}^{n}c_{j}\phi_{j}(t)$ be a polynomial of degree $n$ expressed in the discrete orthonormal basis $\{\phi_{j}\}_{j=0}^{n}$ generated by the Arnoldi process. Suppose $p(t)$ is evaluated at any $t\in[-1,1]$ using the recurrence relation encoded in the upper Hessenberg matrix $H\in\mathbb{R}^{(n+1)\times n}$ in floating-point arithmetic with machine epsilon $\varepsilon_{\text{mach}}$. The computed value, denoted as $\text{fl}(p(t))$, satisfies the forward error bound:
@@
To evaluate the full polynomial $p(t)$, the total accumulated error is the sum of the local errors weighted by the coefficients $c_{j}$. Applying the Cauchy-Schwarz inequality to the inner product of the coefficients and the accumulated basis errors yields:
@@
Let $f\in C^{2k+2}([0,1],\mathbb{R}^{(n-p)\times p})$ be the true coordinate function. Suppose $\hat{f}$ is the global Hermite interpolant of degree $2k+1$ constructed via the CV+A method over $m=k+1$ Chebyshev nodes of the first kind on $[0,1]$. Then the infinity-norm error satisfies:
@@
To rigorously establish the error bound, we decompose the total interpolation error into the analytical truncation error in infinite precision and the numerical round-off error in finite-precision arithmetic. Let $p^{*}(t)$ be the exact theoretical Hermite interpolating polynomial of degree $2k+1$, and let $\hat{f}(t)$ be the numerical interpolant computed via the CV+A algorithm. By the triangle inequality:
@@
For the analytical error, applying the Cauchy remainder theorem for Hermite interpolation component-wise to $f(t)$ yields:
$f(t)-p^{*}(t)=\frac{f^{(2k+2)}(\xi)}{(2k+2)!}\prod_{i=0}^{k}(t-t_{i})^{2},\quad\xi\in(0,1).$
For Chebyshev nodes of the first kind affinely mapped to the interval $[0,1]$, the nodal polynomial $\omega(t)=\prod_{i=0}^{k}(t-t_{i})$ satisfies the classical minimax bound $\max_{t\in[0,1]}|\omega(t)|\leq\frac{1}{2^{2k+1}}$. Since Hermite interpolation matches both values and derivatives, the nodal polynomial is squared, yielding:
@@ 5.3 Total Error Bound
Let $\mathcal{U}(t)$ be a smooth subspace trajectory on the Grassmann manifold $\operatorname{Gr}(n,p)$ for $t\in[0,1]$, and let its MV-coordinate representation $\Xi(t)\in\mathbb{R}^{(n-p)\times p}$ be $C^{2k+2}$ continuous. If $\hat{P}(t)$ is the orthogonal projector reconstructed from the global degree $2k+1$ CV+A Hermite interpolant over $k+1$ Chebyshev nodes, the total subspace error satisfies:
@@
From the geometric error propagation analysis in Theorem 1, the discrepancy between the exact projector $P(t)=\Psi(\Xi(t))$ and the numerically reconstructed projector $\hat{P}(t)=\Psi(\hat{\Xi}(t))$ is bounded by:
@@
Since the reference point for the Stage I Householder stabilization is strategically chosen within the interpolation interval, and $\Xi(t)$ is smooth on the compact set $[0,1]$, the norm of the exact coordinate is globally bounded; that is, $\max_{t\in[0,1]}\|\Xi(t)\|_{2}\leq M$ for some finite constant $M$. As Theorem 3 establishes that the coordinate-domain error for the global degree $2k+1$ CV+A Hermite interpolant satisfies:
@@
Substituting (24) into (23), we get the composition of the geometric and algebraic errors. Because the polynomial degree $k$ drives spectral convergence, the analytical truncation term decays super-exponentially. Consequently, for sufficiently high degrees $k$, the coordinate error shrinks rapidly ($\|\Delta(t)\|_{F}\ll 1$), ensuring that the strictly quadratic term $C_{3}\|\Delta(t)\|_{F}^{2}$ becomes entirely negligible compared to the linear terms. Factoring out the dominant linear contribution and utilizing the intrinsic property $\|\widetilde{U}_{1}\|_{2}\leq 1$ from Theorem 1, gives:
@@
Based on the above analysis, we can see that the Stage I geometric stabilization enforces $\kappa_{\mathrm{geo}}\approx\sqrt{p}$ and ensures the local coordinates remain centered (i.e., $M$ is tightly controlled). Thus, the structural distortion multiplier $(\sqrt{2}+C_{2}M)$ is bounded and strictly independent of the polynomial degree $k$. Absorbing it into a unified geometric projection constant $C$, we obtain the spectral error bound:
@@ 6 Numerical Examples
We consider the Grassmann manifold $\operatorname{Gr}(1000,10)$ with a highly nonlinear curve generated by transcendental functions:
@@
$t_{i}=\frac{1}{2}-\frac{1}{2}\cos\left(\frac{\pi i}{7}\right),\quad i=0,\ldots,7.$ (27)
@@
The Helmholtz equation is known to be difficult to parameterize globally due to physical resonances [11], which cause the solution manifold to oscillate and twist highly nonlinearly with respect to the wavenumber. In this example, we consider the one-dimensional interior Helmholtz boundary value problem parameterized by the wavenumber $k$. The governing equation on the spatial domain $\Omega=[0,1]$ is given by:
@@
Applying a standard spatial discretization (e.g., finite differences) on a uniform grid transforms the continuous problem into a parameter-dependent discrete linear system:

### 27 Mar 2026 v2

@@ 5.2 Algebraic Stability of CV+A
Let $p(t)=\sum_{j=0}^{n}c_{j}\phi_{j}(t)$ be a polynomial of degree $n$ expressed in the discrete orthonormal basis $\{\phi_{j}\}_{j=0}^{n}$ generated by the Arnoldi process. Suppose $p(t)$ is evaluated at any $t\in[-1,1]$ using the recurrence relation encoded in the upper Hessenberg matrix $H\in\mathbb{R}^{(n+1)\times n}$ in floating-point arithmetic with machine epsilon $\varepsilon_{\text{mach}}$. The computed value, denoted as $\text{fl}(p(t))$, satisfies the forward error bound
@@
To evaluate the full polynomial $p(t)$, the total accumulated error is the sum of the local errors weighted by the coefficients $c_{j}$. Applying the Cauchy-Schwarz inequality to the inner product of the coefficients and the accumulated basis errors yields
@@
Let $f\in C^{2k+2}([0,1],\mathbb{R}^{(n-p)\times p})$ be the true coordinate function. Suppose $\hat{f}$ is the global Hermite interpolant of degree $2k+1$ constructed via the CV+A method over $m=k+1$ Chebyshev nodes of the first kind on $[0,1]$. Then the infinity-norm error satisfies
@@
To establish the error bound, we decompose the total interpolation error into the analytical truncation error in infinite precision and the numerical round-off error in finite-precision arithmetic. Let $p^{*}(t)$ be the exact theoretical Hermite interpolating polynomial of degree $2k+1$, and let $\hat{f}(t)$ be the numerical interpolant computed via the CV+A algorithm. By the triangle inequality,
@@
For the analytical error, applying the classical Cauchy remainder theorem for Hermite interpolation component-wise to $f(t)$ and taking the infinity norm on both sides gives the strict upper bound
$\|f(t)-p^{*}(t)\|_{\infty}\leq\frac{\max_{\xi\in[0,1]}\|f^{(2k+2)}(\xi)\|_{\infty}}{(2k+2)!}\prod_{i=0}^{k}(t-t_{i})^{2}.$
For Chebyshev nodes of the first kind affinely mapped from [1,1][-1,1] to the interval $[0,1]$, the monic nodal polynomial $\omega(t)=\prod_{i=0}^{k}(t-t_{i})$ satisfies the classical minimax bound $\max_{t\in[0,1]}|\omega(t)|\leq\frac{1}{2^{2k+1}}$. Since Hermite interpolation matches both values and first derivatives, the nodal polynomial is squared, yielding the global analytical bound
@@ 5.3 Total Error Bound
Let $\mathcal{U}(t)$ be a smooth subspace trajectory on the Grassmann manifold $\operatorname{Gr}(n,p)$ for $t\in[0,1]$, and let its MV-coordinate representation $\Xi(t)\in\mathbb{R}^{(n-p)\times p}$ be $C^{2k+2}$ continuous. If $\hat{P}(t)$ is the orthogonal projector reconstructed from the global degree $2k+1$ CV+A Hermite interpolant over $k+1$ Chebyshev nodes, the total subspace error satisfies
@@
From the geometric error propagation analysis in Theorem 1, the discrepancy between the exact projector $P(t)=\Psi(\Xi(t))$ and the numerically reconstructed projector $\hat{P}(t)=\Psi(\hat{\Xi}(t))$ is bounded by
@@
Since the reference point for the Stage I Householder stabilization is strategically chosen within the interpolation interval, and $\Xi(t)$ is smooth on the compact set $[0,1]$, the norm of the exact coordinate is globally bounded; that is, $\max_{t\in[0,1]}\|\Xi(t)\|_{2}\leq M$ for some finite constant $M$. As Theorem 3 establishes that the coordinate-domain error for the global degree $2k+1$ CV+A Hermite interpolant satisfies
@@
Substituting (24) into (23), we get the composition of the geometric and algebraic errors. Because the polynomial degree $k$ drives spectral convergence, the analytical truncation term decays super-exponentially. Consequently, for sufficiently high degrees $k$, the coordinate error shrinks rapidly ($\|\Delta(t)\|_{F}\ll 1$), ensuring that the strictly quadratic term $C_{3}\|\Delta(t)\|_{F}^{2}$ becomes entirely negligible compared to the linear terms. Factoring out the dominant linear contribution and utilizing the intrinsic property $\|\widetilde{U}_{1}\|_{2}\leq 1$ from Theorem 1, gives
@@
Based on the above analysis, we can see that the Stage I geometric stabilization enforces $\kappa_{\mathrm{geo}}\approx\sqrt{p}$ and ensures the local coordinates remain centered (i.e., $M$ is tightly controlled). Thus, the structural distortion multiplier $(\sqrt{2}+C_{2}M)$ is bounded and strictly independent of the polynomial degree $k$. Absorbing it into a unified geometric projection constant $C$, we obtain the spectral error bound
@@ 6 Numerical Examples
We consider the Grassmann manifold $\operatorname{Gr}(1000,10)$ with a highly nonlinear curve generated by transcendental functions
@@
$t_{i}=\frac{1}{2}-\frac{1}{2}\cos\left(\frac{\pi i}{7}\right),\quad i=0,\ldots,7,$ (27)
@@
The Helmholtz equation is known to be difficult to parameterize globally due to physical resonances [11], which cause the solution manifold to oscillate and twist highly nonlinearly with respect to the wavenumber. In this example, we consider the one-dimensional interior Helmholtz boundary value problem parameterized by the wavenumber $k$. The governing equation on the spatial domain $\Omega=[0,1]$ is given by
@@
Applying a standard spatial discretization (e.g., finite differences) on a uniform grid transforms the continuous problem into a parameter-dependent discrete linear system